OpenStudy (anonymous):

how to solve for r in S=a/1-r

OpenStudy (anonymous):

r=a-s

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

book says r=S-a/S

OpenStudy (anonymous):

r=a-s s=a-r a=r+s

OpenStudy (anonymous):

do u get it

OpenStudy (anonymous):

what is your answer choices

OpenStudy (anonymous):

there are no answer choices just the answer key which say the answer is r=S-a/S but does not show the steps to solve it

OpenStudy (anonymous):

it is a no solution

OpenStudy (anonymous):

i got it! thank you (:

OpenStudy (anonymous):

what was it

OpenStudy (anonymous):

r=S-a/S I was skipping a step